OpenStudy (anonymous):

Explain, in complete sentences, how to factor 4x2 – 4x – 48. Explain how your factors could be checked for accuracy

OpenStudy (anonymous):

let f(x)= 4x^2 – 4x – 48=4(x^2-x-12) f(4)=0 therefore, x-4 is a factor of f(x) f(-3)=0 therefore, x+3 is a factor of f(x) deg(f(x))=2 therefore, it has two roots therefore, f(x) =4(x-4)(x+3)

OpenStudy (radar):

Extract the common factor 4 getting 4(x2-x-12) then factor the remainder getting: (x-4)(x+3) put them together and get 4(x-4)(x+3)

OpenStudy (radar):

To check for accuracy simply multiply your results to determine if you end up with the original expression

OpenStudy (anonymous):

The first step to factorization is to take out a highest common factor, which in your case is 4 . 4(x^2-x-12) Now you have to ask yourself what multiplys to give -12 and and adds to give -1, well its right, -4,3 4(x-4)(x+3) to check you can expand again and you should end up with a original question

OpenStudy (anonymous):

Thank you ,this helps alot